Henry invested $12,000 in a new restrauant. When the restaurant was sold two years later, he recieved $27,000.Find his average annual return by solving the equatin 12,000(1+r)^2=27,000.
Answer by unlockmath(1688)

Hello,
The key to finding the answer is getting r all alone in this: 12,000(1+r)^2=27,000
This can be rewritten as:
12,000(r^2+2r+1)=27,000 and rewritten again as:
12,000r^2+24,000r+12,000=27,000 Now let's subtract 27,000 from both sides to give us:
12,000r^2+24,000r-15,000=0 Now take out 3000 to be:
3000(4r^2+8r-5)=0 Now use the quadratic formula to solve x:
{{r = (-b +- sqrt( b^2-4*a*c ))/(2*a) }} Plugging in the values we get:
{{r = (-8 +- sqrt( 8^2+80))/(8) }}This works out to .5 the other value is negative so we have:
rate = .5 or 50% Good return for that investment!
This can be checked by plugging the answer into the original equation.
RJ
Check out a new book I wrote at:
www.math-unlock.com